# Lyman Continuum leakage from massive leaky starbursts: A different class of emitters?

### v1

@@ Abstract
The origin of Lyman continuum (LyC) photons responsible for reionizing the universe remains a mystery, with the fraction of escaping LyC photons from galaxies at 6 to 12 being highly uncertain. While direct detection of LyC photons from this epoch is hindered by absorption from the intergalactic medium, lower-redshift analogs offer a promising avenue to study LyC leakage. We present Hubble Space Telescope Cosmic Origins Spectrograph (HST COS) observations of five low-redshift (z 0.3) massive starburst galaxies, selected for their high stellar mass and weak [S II] nebular emission—an indirect tracer of LyC escape. Three of the five galaxies show LyC leakage, highlighting the reliability of weak [S II] as a tracer, especially in light of recent JWST discoveries of z 5 galaxies with similarly weak [S II] emission. The dust-corrected LyC escape fractions (), which represent the LyC photons that would escape in the absence of dust, range from 33% to 84%. However, the absolute escape fractions (), which show the LyC photons escaping after passing through both neutral hydrogen HI absorption and dust attenuation, are significantly lower, ranging between 1% and 3%. This suggests that while the galaxies are nearly optically thin to HI, their high dust content significantly suppresses LyC photon escape. These [SII]-weak, massive leakers are distinct from typical low-redshift LyC emitters, showing higher metallicity, lower ionization states, more dust extinction and higher star formation surface densities. This suggests that these galaxies constitute a distinct population, likely governed by a different mechanism facilitating LyC photon escape. We propose that the feedback-driven winds in these compact starbursts create ionized channels through which LyC photons escape, aligning with a “picket-fence” model.
@@ 1 Introduction
Although many attempts have been made to observationally constrain the escape fraction in star forming galaxies across a wide range of redshift, most of the early observations yielded upper limits only with a few detections, indicating 3% (Leitherer et al., 1995; Deharveng et al., 2001; Malkan et al., 2003; Alavi et al., 2020). With a more extensive search program in recent years using the Hubble Space Telescope (HST) Cosmic Origins Spectrograph (COS; Green et al., 2012), convincing detections of LyC leaking emission have been found in a few tens of sources, mostly from local (0.2z0.4) starbursts, “Green Pea” galaxies, and Lyman break analogs (Heckman et al., 2005; Hoopes et al., 2007; Iwata et al., 2009; Borthakur et al., 2014; Izotov et al., 2016b, a, 2018a, 2018b, 2021; Wang et al., 2019, 2021; Flury et al., 2022a; Xu et al., 2022). Notably, the recently completed Low Redshift Lyman Continuum Survey (LzLCS; Saldana-Lopez et al., 2022; Flury et al., 2022a, b) assembled a sample of 66 star forming LyC leaking “candidate” galaxies from the Sloan Digital sky survey (SDSS: York et al., 2000) and Galaxy Evolution Explorer (GALEX: Martin et al., 2005), which were observed with HST/COS. The program added 35 new LyC detections with 97.7% confidence. These Lyman continuum emitters (LCEs) serve as the largest Lyman Continuum Emitters sample detected to date.
@@
In contrast, Borthakur et al. (2014) and Wang et al. (2019) found four galaxies which appear to be an entirely different class of LyC leaking sources. These galaxies exhibit properties entirely different from the traditionally known LCEs and are characterized by large stellar mass , a much lower ionization state ([OIII]/[OII] 1), with the UV emission dominated by an extremely compact (few hundred parsecs) central region. They also display a relative weakness of [SII]6717Å, 6731 Å lines defined with respect to typical star forming galaxies. It has been argued that a different mechanism may be at play here: these massive and compact starbursts drive extreme galactic winds, which blow holes and clear out low density channels in the neutral ISM through which LyC photons can escape (e.g., Bergvall et al., 2006; Heckman et al., 2011). In a classical HII region, a [SII] deficiency indicates a depleted partially ionized region near the outer edge of the Stromgren sphere (Pellegrini et al., 2012; Borthakur et al., 2014; Heckman et al., 2011; Alexandroff et al., 2015). In the massive, compact leaky galaxies this is possibly due to the fully ionized channels created by the feedback driven outflows.
Thus, there may be two different mechanisms to create channels in the ISM to enable LyC escape. In this study, we expand the second sample of members of the massive, compact leaky starbursts by observing five new candidates using HST/COS observations. These galaxies are selected to be [SII]-deficient, massive star forming galaxies (selection criteria further described in §2). We use COS G140L observations that reach below the Lyman Limit to directly measure the amount of escaping LyC radiation, and to derive the LyC escape fraction. We also study the LyC escape fraction and the nature of correlations with host galaxy properties, and compare the properties of this population of galaxies with the traditional LCEs from the LzLCS survey (Flury et al., 2022a, b; Saldana-Lopez et al., 2022). Our main goal is to determine whether there exist a real dichotomy in the properties of the leaky galaxies in these two classes, and how this may depend on the physical condition of the host galaxy and its ISM.
@@ 2 Sample Selection
The COS NUV acquisition images of the sources are shown in Fig. 1. The targets and their relevant properties are listed in Table. 1.
ID | z | [OIII]/[OII] | H EW | Ly EW | 12 + log (O/H) | ||||||||
[Å] | [Å] | [] | [] | [ | |||||||||
J0121 | 0.264 | 0.17 | 0.713 | 0.0043 | 0.515 | 21 | 43 | 0.298 | 8.59 | 10.5 | 60 | 130.9 | -0.278 |
J0219 | 0.327 | 0.017 | 0.05 | 0.0035 | 0.52 | 13 | 8 | 0.235 | 8.48 | 10.5 | 26 | 22.1 | -0.457 |
J1145 | 0.287 | 0.221 | 0.841 | 0.0264 | 0.490 | 26 | 1 | 0.345 | 8.48 | 10.0 | 25 | 6.7 | -0.249 |
J1230 | 0.342 | 0.078 | 0.333 | 0.011 | 0.437 | 18 | 11 | 0.30 | 8.49 | 10.4 | 36 | 13.5 | -0.213 |
J2348 | 0.313 | 0.022 | 0.14 | 0.0025 | 0.607 | 16 | 32 | 0.351 | 8.53 | 10.6 | 53 | 31.4 | -0.169 |
Throughout the study, we use the term "massive leakers" to refer to the combined sample of 5 sources from this work, the three massive, [SII]-deficit LyC leakers previously studied in Wang et al. (2019) and the massive LyC leaker from Borthakur et al. (2014). We compare these massive leaker population with the more traditional LCE sample from the low-z Lyman continuum leaker survey (LzLCS: Flury et al., 2022a; Saldana-Lopez et al., 2022). The LzLCS sample includes 66 candidate low redshift (z 0.3-0.4) leakers selected from SDSS & GALEX data archive. The sample spans a broad range in host galaxy parameters like [OIII]/[OII], , stellar mass, metallicity, UV slope and others. The program was carried on with 134 orbits of HST/ COS spectroscopy a part of GO 15626 (Cycle 26, PI: A. Jaskot). All the measurements of host galaxy properties of the LzLCS sample are obtained from the catalog presented in Flury et al. (2022a, b). In addition to the 66 LzLCS galaxies, the catalog also compiles the measurements of 23 other LCE candidates from the literature observed by HST/COS as part of other programs (Izotov et al., 2016b, a, 2018a, 2018b, 2021). The HST/COS observations were reprocessed and the LyC fluxes were recomputed systematically by Flury et al. (2022a) to obtain consistent LyC measurements. We will refer to this combined sample of 89 galaxies as our comparison sample, and often loosely as the LzLCS+ sample (since it includes other LCEs in addition to the LzLCS survey).
@@ 3.2 Measuring Lyman Continuum
We smooth the spectra of all sources with a Gaussian kernel with full width at half maximum of about 1.0 Å, to achieve the native resolution and boost the relatively low signal-to-noise. We measure the LyC flux in each source in a rest frame 20 Å window as close as possible to = 900 Å, while simultaneously avoiding contamination from Ly and NI 1200 geocoronal emission lines. Thus, we avoid wavelengths above = 1180 Å (similar to Flury et al., 2022a). We also do not choose the spectral range to include wavelengths below = 1100 Å, since the sensitivity of the G140L grating and the signal-to-noise of the spectra decline rapidly. Thus, to compute LyC flux () for our sources at z = 0.26 0.34, we compute the mean background subtracted flux density in the chosen 20 Å spectral window generally within = 860-910 Å, such that the above conditions are met. This ensures a uniform measurement of the LyC flux across our entire sample and is also consistent with the LzLCS survey to enable direct comparison.
One of the main goals of this investigation is to explore the roles of both dust extinction and absorption by HI in determining the fraction of escaping LyC photons. We also want a mix of indicators that includes a simple model-independent parameter. We will therefore characterize the escape of LyC photons by reporting three different quantities: 1) The direct flux ratio /Fλ1100 (often labelled as ) 2) the ratio of FλLyC to the intrinsic contribution from the stars, as inferred from fitting models of the observed UV continuum (often labelled as absolute escape fraction ), and 3) (often labelled as ), which is corrected for the effects of dust attenuation and represents the fraction of LyC photons that would escape after surviving neutral hydrogen absorption, assuming no dust is present.
To derive /Fλ1100, we transform the observed spectra to the rest frame of the galaxy using SDSS spectroscopic redshifts. We compute by calculating the mean flux density in the LyC spectral window of width 20 Å, as discussed above. Fλ1100 is calculated by taking average flux in 1090 1100 Å, i.e. keeping the spectral window width to be the same as the LyC flux for consistency. Our method is consistent with previous studies, allowing for direct comparison (Wang et al., 2019; Flury et al., 2022a).
@@
To perform the fitting, the observed spectra are first manually placed into the rest-frame by multiplying by the corresponding 1/(1 + z) factor. Both the spectra and the models were then normalized by the median flux within a wavelength interval free of stellar and ISM features (1350–1370 Å), and all the fits are performed in the same rest-wavelength range (930–1425 Å). Finally, the models are convolved by a Gaussian kernel to the instrumental resolution (R 1000). The best-fit stellar model spectrum for each source is integrated over the optimal 20 Å wide LyC spectral window to estimate the modeled (i.e. intrinsic) LyC flux. Fig. 3 shows the FICUS fitting for one of the galaxies (ID: J0121, at z = 0.264), with the observed spectrum in black and fitted stellar continuum in red.
@@
We use FICUS for stellar continuum modeling to be consistent with the results from the LzLCS+ study. However, for one of our sources (ID: J1145), the FICUS fit fails due to inadequate signal to noise, particularly at the red end of the spectrum. So for J1145, we directly generate a grid of synthetic spectra from SB99 models, and assume a continuous and constant star formation rate (SFR) = 1 , and a Kroupa Initial mass function (Kroupa, 2001). We generate a total of eight sets of SB99 models based on two choices each for burst age ( and yr), metallicity (solar or 1/7 solar), and whether or not stellar rotation is present. We assume the stellar population evolves from the zero-age main sequence using the evolutionary models of the Geneva Group (similar to the FICUS assumptions). The model spectra are interpolated into the same wavelength array as its corresponding COS spectrum, convolved with the same Gaussian kernel, and normalized to the common flux level (1350–1370 Å) – following the same procedure as the rest of the sources. The best fit is chosen by eye; more specifically, we closely examine the match of the two strong stellar wind features – O VI 1032,1038 and N V 1238,1242 (similar to Wang et al., 2019). These P-Cygni features trace the most massive stars that are responsible for producing most of the ionizing continuum. We multiply the SB99 model-derived LyC flux with the total SFR since the stellar models were assumed to have a constant SFR = 1 in the model grid. The SFR was derived using WISE band 3 and band 4 data to estimate a rest-frame 24 micron luminosity and then using the relationship in Kennicutt & Evans (2012) to get SFRIR. We can then use this SFR and the SB99 model to obtain the intrinsic LyC flux ().
@@
Then is essentially the ratio of / . The measured value of this escape fraction then depends on both the effects of dust extinction and photoelectric absorption of the LyC due to hydrogen.
@@
For method 3, we measure the dust-corrected escape fraction which is the ratio of emergent LyC flux after removing the effects of dust extinction compared to the intrinsic flux. To determine the amount of dust extinction, we first measure the rest-frame 1500 Å flux and convert that to luminosity. We use the to estimate using the Kennicutt & Evans (2012) relation, and derive the attenuation at 1500 Å by:
@@
The far-UV dust attenuation law from Reddy et al. (2016) suggests that E(B-V)900 = 0.226 A1500, extrapolated from 950 to 900 Å based on their fit to data between 950 and 2000 Å. The dust-corrected ratio is calculated using . The final estimate is obtained by dividing the dust-corrected by the intrinsic derived from the modeled spectrum. This value measures only the photoelectric absorption of LyC photons by HI in the absence of dust, providing complementary information to . Throughout the text, is also referred to as .

### v2

@@ Abstract
The origin of Lyman continuum (LyC) photons responsible for reionizing the universe remains largely unknown, with the fraction of escaping LyC photons from galaxies at to 12 still uncertain. Direct detection of LyC photons from this epoch is challenging due to intergalactic medium absorption, making lower-redshift analogs valuable for studying LyC leakage. In this study, we present Hubble Space Telescope Cosmic Origins Spectrograph (HST COS) observations of five low-redshift (z 0.3) massive starburst galaxies, selected for high stellar mass and weak [S II] nebular emission, an indirect tracer of LyC escape. LyC leakage is detected in three of the five galaxies, highlighting weak [S II] as a reliable tracer — a finding supported by recent JWST discoveries of z 5 galaxies with similarly weak [S II] emission. The dust-corrected LyC escape fractions (), representing LyC photons that would escape without dust, range from 33% to 84%. However, the absolute escape fractions (), accounting for both neutral hydrogen absorption and dust attenuation, are substantially lower, between 1% and 3%. This indicates that, although these galaxies are nearly optically thin to HI, their significant dust content restricts LyC escape. These [S II]-weak, massive leakers differ from typical low-redshift LyC emitters, exhibiting higher metallicity, lower ionization states, greater dust extinction, and higher star formation surface densities. We suggest that feedback-driven winds in these compact starbursts generate ionized channels, allowing LyC escape in line with a "picket-fence" model, indicating a distinct mechanism for LyC leakage.
@@ 1 Introduction
Although many attempts have been made to observationally constrain the escape fraction in star forming galaxies across a wide range of redshift, most of the early observations yielded upper limits only with a few detections, indicating 3% (Leitherer et al., 1995; Deharveng et al., 2001; Malkan et al., 2003; Alavi et al., 2020). With a more extensive search program in recent years using the Hubble Space Telescope (HST) Cosmic Origins Spectrograph (COS; Green et al., 2012), convincing detections of LyC leaking emission have been found in a few tens of sources, mostly from local (0.2z0.4) starbursts, “Green Pea” galaxies, and Lyman break analogs (Heckman et al., 2005; Hoopes et al., 2007; Iwata et al., 2009; Borthakur et al., 2014; Izotov et al., 2016b, a, 2018a, 2018b, 2021; Wang et al., 2019, 2021; Flury et al., 2022a; Xu et al., 2022). Notably, the recently completed Low Redshift Lyman Continuum Survey (LzLCS+; Saldana-Lopez et al., 2022; Flury et al., 2022a, b) assembled a sample of 66 star forming LyC leaking “candidate” galaxies from the Sloan Digital sky survey (SDSS: York et al., 2000) and Galaxy Evolution Explorer (GALEX: Martin et al., 2005), which were observed with HST/COS. The program added 35 new LyC detections with 97.7% confidence. These Lyman continuum emitters (LCEs) serve as the largest Lyman Continuum Emitters sample detected to date. In addition, a few Lyman continuum detections with high escape fractions have also been reported at high redshifts (Vanzella et al., 2018; Fletcher et al., 2019; Vanzella et al., 2019).
@@
In contrast, Borthakur et al. (2014) and Wang et al. (2019) found four galaxies which appear to be an entirely different class of LyC leaking sources. These galaxies exhibit properties entirely different from the traditionally known LCEs and are characterized by large stellar mass , a much lower ionization state ([OIII]/[OII] 1), with the UV emission dominated by an extremely compact (few hundred parsecs) central region. They also display a relative weakness of [SII]6717Å, 6731 Å lines defined with respect to typical star forming galaxies. A deficit in [S II] emission has been proposed as an indicator of galaxies that are optically thin to ionizing radiation. Recent JWST observations have revealed several high-redshift () galaxies with extremely weak [S II] lines (Sanders et al., 2023; Cameron et al., 2023; Hayes et al., 2025). In particular, Hayes et al. (2025) show that the most highly ionized high- galaxies, when stacked, frequently display a pronounced [S II] deficit and emerge as outliers in multiple properties, including slope, Mg II escape fraction, and ionizing photon production efficiency. These characteristics closely resemble those of the [S II]-deficient Lyman continuum leakers reported by Wang et al. (2019).
It has been argued that a different mechanism may be at play here: these massive and compact starbursts drive extreme galactic winds, which blow holes and clear out low density channels in the neutral ISM through which LyC photons can escape (e.g., Bergvall et al., 2006; Heckman et al., 2011). In a classical HII region, a [SII] deficiency indicates a depleted partially ionized region near the outer edge of the Stromgren sphere (Pellegrini et al., 2012; Borthakur et al., 2014; Heckman et al., 2011; Alexandroff et al., 2015). In the massive, compact leaky galaxies this is possibly due to the fully ionized channels created by the feedback driven outflows.
Thus, there may be two different mechanisms to create channels in the ISM to enable LyC escape. A third possibility has also been proposed, in which merger-driven interactions disrupt the ISM and produce such low-density channels (Le Reste et al., 2024; Yuan et al., 2021; Zhu et al., 2024; Mascia et al., 2024). However, we do not explore this scenario here, as detailed morphological analyses and merger identification for our sources lie beyond the scope of this work. In this study, we expand the second sample of members of the massive, compact leaky starbursts by observing five new candidates using HST/COS observations. These galaxies are selected to be [SII]-deficient, massive star forming galaxies (selection criteria further described in §2). We use COS G140L observations that reach below the Lyman Limit to directly measure the amount of escaping LyC radiation, and to derive the LyC escape fraction. We also study the LyC escape fraction and the nature of correlations with host galaxy properties, and compare the properties of this population of galaxies with the traditional LCEs from the LzLCS+ survey (Flury et al., 2022a, b; Saldana-Lopez et al., 2022). Our main goal is to determine whether there exist a real dichotomy in the properties of the leaky galaxies in these two classes, and how this may depend on the physical condition of the host galaxy and its ISM.
@@ 2 Sample Selection
The COS NUV acquisition images of the sources are shown in Fig. 1. The targets and their relevant properties are listed in Table. 1 and 2. All the HST data used in this work can be downloaded from MAST using the DOI: 10.17909/63mb-ag19.
| ID | z | Signif. | ||||||
| [/Å] | ||||||||
| J0121 | 0.264 | 2.90.9 | 2.6 | 0.170.05 | 0.7130.23 | 0.00430.0013 | 0.00590.0028 | |
| J0219* | 0.327 | 0.480.93 | 0.0170.03 | 0.050.1 | 0.00350.007 | 0.070.198 | ||
| J1145 | 0.287 | 5.381.10 | 5.7 | 0.2210.04 | 0.8410.16 | 0.02640.0052 | 0.03140.0087 | |
| J1230 | 0.342 | 1.580.51 | 2.8 | 0.0780.02 | 0.3330.11 | 0.0110.004 | 0.03303 0.0169 | |
| J2348* | 0.313 | 0.300.82 | 2 | 0.0220.06 | 0.140.30 | 0.00250.005 | 0.0178 0.0505 |
* marked sources represent Lyman Continuum non emitters and the reported flux and escape fractions are upper limits, derived as described in §4.1.
Throughout the study, we use the term "massive leakers" to refer to the combined sample of five sources from this work, the three massive, [SII]-deficit LyC leakers previously studied in Wang et al. (2019) and the massive LyC leaker from Borthakur et al. (2014). We compare these massive leaker population with the more traditional LCE sample from the low-z Lyman continuum leaker survey (LzLCS+: Flury et al., 2022a; Saldana-Lopez et al., 2022). The LzLCS+ sample includes 66 candidate low redshift (z 0.3-0.4) leakers selected from SDSS & GALEX data archive. The sample spans a broad range in host galaxy parameters like [OIII]/[OII], star formation rate surface density (), stellar mass, metallicity, UV slope and others. The program was carried on with 134 orbits of HST/ COS spectroscopy a part of GO 15626 (Cycle 26, PI: A. Jaskot). All the measurements of host galaxy properties of the LzLCS+ sample are obtained from the catalog presented in Flury et al. (2022a, b). In addition to the 66 galaxies, the catalog also compiles the measurements of 23 other LCE candidates from the literature observed by HST/COS as part of other programs (Izotov et al., 2016b, a, 2018a, 2018b, 2021). The HST/COS observations were reprocessed and the LyC fluxes were recomputed systematically by Flury et al. (2022a) to obtain consistent LyC measurements. We will refer to this combined sample of 89 galaxies as our LzLCS+ sample.
@@ 3.2 Measuring Lyman Continuum
We smooth the spectra of all sources with a Gaussian kernel with full width at half maximum of about 1.0 Å, to achieve the native resolution and boost the relatively low signal-to-noise. We measure the LyC flux in each source in a rest frame 20 Å window as close as possible to = 900 Å, while simultaneously avoiding contamination from Ly and NI 1200 geocoronal emission lines. Thus, we avoid wavelengths above = 1180 Å (similar to Flury et al., 2022a). We also do not choose the spectral range to include wavelengths below = 1100 Å, since the sensitivity of the G140L grating and the signal-to-noise of the spectra decline rapidly. Thus, to compute LyC flux () for our sources at z = 0.26 0.34, we compute the mean background subtracted flux density in the chosen 20 Å spectral window generally within = 860-910 Å, such that the above conditions are met. This ensures a uniform measurement of the LyC flux across our entire sample and is also consistent with the LzLCS+ survey to enable direct comparison.
To robustly assess the LyC detections, we calculate the probability, , that the total observed (gross) counts within the LyC spectral window arise from the distribution of background counts , following the definition of Worseck et al. (2016) and the methodology adopted by Flury et al. (2022a). We employ the Python package FeldCous (Flury, 2025; Feldman & Cousins, 1998) to compute and derive the detection significance, using the same prescription as Flury et al. (2022a), to identify galaxies that meet the conventional detection threshold. The resulting detection significances for our sources are listed in Table 1.
One of the main goals of this investigation is to explore the roles of both dust extinction and absorption by HI in determining the fraction of escaping LyC photons. We also want a mix of indicators that includes a simple model-independent parameter. We will therefore characterize the escape of LyC photons by reporting three different quantities:
The direct flux ratio /Fλ1100 (often labelled as ) which we denote as Method 1.
, which is the ratio of FλLyC to the intrinsic contribution from the stars, as inferred from fitting models of the observed UV continuum (often labelled as absolute escape fraction ). This is Method 2.
(often labelled as ), which is corrected for the effects of dust attenuation and represents the fraction of LyC photons that would escape after surviving neutral hydrogen absorption, assuming no dust is present. This is denoted by Method 3.
To derive /Fλ1100, we transform the observed spectra to the rest frame of the galaxy using SDSS spectroscopic redshifts. We compute by calculating the mean flux density in the LyC spectral window of width 20 Å, as discussed above. Fλ1100 is calculated by taking average flux in 1090 1100 Å, i.e. keeping the spectral window width to be the same as for consistency. Our method is consistent with previous studies, allowing for direct comparison (Wang et al., 2019; Flury et al., 2022a).
@@
To perform the fitting, the observed spectra are first manually placed into the rest-frame by multiplying by the corresponding 1/(1 + z) factor. Both the spectra and the models are then normalized by the median flux within a wavelength interval free of stellar and ISM features (1350–1370 Å), and all the fits are performed in the same rest-wavelength range (930–1425 Å). Finally, the models are convolved by a Gaussian kernel to the instrumental resolution (R 1000). The best-fit stellar model spectrum for each source is integrated over the optimal 20 Å wide LyC spectral window to estimate the modeled (i.e. intrinsic) LyC flux (). Fig. 3 shows the FICUS fitting for one of the galaxies (ID: J0121, at z = 0.264), with the observed spectrum in black and fitted stellar continuum in red.
@@
We use FICUS for stellar continuum modeling to be consistent with the results from the LzLCS+ study. However, for one of our sources (ID: J1145), the FICUS fit fails due to inadequate signal to noise, particularly at the red end of the spectrum. So for J1145, we directly generate a grid of synthetic spectra from STARBURST99 models, and assume a continuous and constant star formation rate (SFR) = 1 , and a Kroupa Initial mass function (Kroupa, 2001). We generate a total of eight sets of SB99 models based on two choices each for burst age ( and yr), metallicity (solar or 1/7 solar), and whether or not stellar rotation is present. We assume the stellar population evolves from the zero-age main sequence using the evolutionary models of the Geneva Group (similar to the FICUS assumptions). The model spectra are interpolated into the same wavelength array as its corresponding COS spectrum, convolved with the same Gaussian kernel, and normalized to the common flux level (1350–1370 Å) – following the same procedure as the rest of the sources. The best fit is chosen by eye; more specifically, we closely examine the match of the two strong stellar wind features – O VI 1032,1038 and N V 1238,1242 (similar to Wang et al., 2019). These P-Cygni features trace the most massive stars that are responsible for producing most of the ionizing continuum. We multiply the SB99 model-derived LyC flux with the total SFR since the stellar models were assumed to have a constant SFR = 1 in the model grid. The SFR was derived using WISE band 3 and band 4 data to estimate a rest-frame 24 micron luminosity and then using the relationship in Kennicutt & Evans (2012) to get SFRIR. We then use this SFR and the SB99 model to obtain the intrinsic LyC flux ().
@@
Then is essentially the ratio of / . The measured value of this escape fraction depends on both the effects of dust extinction and photoelectric absorption of the LyC due to hydrogen. The computed values derived from the FICUS package and those obtained directly from Starburst99 differ by less than 1% for all our sources (except J1145 where the FICUS fit fails), confirming a robust characterization of the escape fraction independent of the method used.
@@
For method 3, we measure the dust-corrected escape fraction which is the ratio of emergent LyC flux to the intrinsic flux, after removing the effects of dust extinction. To determine the amount of dust extinction, we first measure the rest-frame 1500 Å flux and convert that to luminosity. We use the to estimate using the Kennicutt & Evans (2012) relation, and derive the attenuation at 1500 Å by:
@@
According to the far-UV dust attenuation law by Reddy et al. (2016), E(-) is given by 0.226 A1500, where = 900 Å and = 1100 Å, based on an extrapolation of their fit to data between 950 and 2000 Å. The dust-corrected ratio is then calculated as . The final estimate is obtained by dividing the dust-corrected by the intrinsic derived from the modeled spectrum. This value measures only the photoelectric absorption of LyC photons by HI in the absence of dust. Thus throughout the text, is also referred to as . Table 1 presents the Lyman Continuum escape fractions, along with their associated uncertainties, calculated using all three methodologies discussed above for our five target objects.
| ID | [OIII]/[OII] | H EW | Ly EW | 12 + log (O/H) | |||||
| [Å] | [Å] | [] | [] | [ | |||||
| J0121 | 0.5150.016 | 21.1 0.4 | 43.13.5 | 0.2980.007 | 8.590.13 | 10.510.1 | 602 | 130.95.2 | -0.278 0.08 |
| J0219 | 0.520.019 | 13.00.5 | 8.02.1 | 0.2350.013 | 8.480.06 | 10.50.11 | 264 | 22.14.1 | -0.457 0.22 |
| J1145 | 0.4900.015 | 25.90.8 | 15 | 0.3450.009 | 8.480.07 | 10.000.03 | 251 | 6.70.7 | -0.249 0.05 |
| J1230 | 0.4370.018 | 18.00.6 | 11.22.5 | 0.3000.015 | 8.490.12 | 10.400.03 | 362 | 13.52.5 | -0.213 0.11 |
| J2348 | 0.6070.02 | 16.20.5 | 32.07.5 | 0.3510.01 | 8.530.13 | 10.600.02 | 532 | 31.43.4 | -0.169 0.03 |